Tricia is 14 years older than Michella. In four years, Tricia will be twice as old as Michella. How old are they now?
Answer by mukhopadhyay(490)

Let Michella's current age is x
=> Tricia's current age is (x+14)
Four years later Michella's age will be (x+4) and Tricia's age will be (x+18);
Per question, (x+18) = 2(x+4)
=> x+18 = 2x+8
=> x = 10
Answer: Michella's current age is 10 years and Tricia's current age is 24 years